OpenStudy (anonymous):

If y = 144 when x = 24, what is y when x = 54?

OpenStudy (igreen):

Set up a proportion: \(\dfrac{144}{24} = \dfrac{y}{54}\) Multiply 54 to both sides: \(y = \dfrac{144 \times 54}{24}\) Simplify

OpenStudy (anonymous):

324?

OpenStudy (igreen):

@cvliboi Let the asker do the math.

OpenStudy (anonymous):

my bad. @iGreen